# Data-driven model order reduction for wave propagation in materials with nonlinearities or damage

### v1

@@ 3.3 The intrusive POD-Galerkin ROM for second order mechanical systems
In this subsection, we describe the intrusive POD-Galerkin ROM for a mechanical system of second order without damping. For the derivation of a reduced order approximation of the full order model (FOM) (12) and (13), respectively, the FOM displacement field $\mathbf{u}(\bm{\theta},t)$ is approximated using a Galerkin ansatz as follows:
$\mathbf{u}(\bm{\theta},t)\approx\mathbf{\Phi}\mathbf{u}_{r}(\bm{\theta},t),$ (18)
with coefficient vector $\mathbf{u}_{r}(\bm{\theta},t)\in\mathbb{R}^{r}$ of dimension $r\ll N$ and reduced basis matrix $\mathbf{\Phi}=[\bm{\phi}_{1},\dots,\bm{\phi}_{r}]\in\mathbb{R}^{N\times r}$. Here, the reduced basis vectors $\{\bm{\phi}_{i}\}_{i=1}^{r}$ are generated using the POD method. For this a SVD of the snapshot matrix of the displacement can be used:
@@
By inserting the reduction assumption (18) in (12) and (13), respectively and performing a Galerkin projection onto the POD space, one may obtain the following linear reduced order dynamical system for a fixed parameter $\bm{\theta}\in\mathcal{D}$:
$\mathbf{\Phi}^{T}\mathbf{M}\mathbf{\Phi}\;\ddot{\mathbf{u}}_{r}(\bm{\theta},t)+\mathbf{\Phi}^{T}\mathbf{K}(\bm{\theta})\mathbf{\Phi}\;\mathbf{u}_{r}(\bm{\theta},t)=\mathbf{\Phi}^{T}\mathbf{f}(t),\quad\textrm{for}\quad t\in[0,T],$ (20)
@@
$\mathbf{\Phi}^{T}\mathbf{M}\mathbf{\Phi}\;\ddot{\mathbf{u}}_{r}(\bm{\theta},t)+\mathbf{g}(\mathbf{\Phi}\;\mathbf{u}_{r}(\bm{\theta},t))=\mathbf{\Phi}^{T}\mathbf{f}(t).$ (21)
@@
$\mathbf{M}_{r}=\mathbf{\Phi}^{T}\mathbf{M}\mathbf{\Phi},\quad\mathbf{K}_{r}(\bm{\theta})=\mathbf{\Phi}^{T}\mathbf{K}(\bm{\theta})\mathbf{\Phi},\quad\mathbf{f}_{r}=\mathbf{\Phi}^{T}\mathbf{f}(t).$ (22)
@@
$\mathbf{M}_{r}\ddot{\mathbf{u}}_{r}(\bm{\theta},t)+\mathbf{K}_{r}(\bm{\theta})\mathbf{u}_{r}(\bm{\theta},t)=\mathbf{f}_{r},\quad\textrm{for}\quad t\in[0,T].$ (23)
@@
Remark 3.1 (Hyper reduction).
We note that the reduced stiffness matrix $\mathbf{K}_{r}(\bm{\theta})$ in (23) depends on the parameter $\bm{\theta}$ and is in general not parameter separable, i.e. in general it cannot be written in the form $\mathbf{K}_{r}(\bm{\theta})=\sum_{q=1}^{Q}\alpha_{q}(\bm{\theta})\mathbf{k}_{q}$ with parameter dependent coefficients $\alpha_{q}(\bm{\theta})\in\mathbb{R}$ and parameter independent matrices $\mathbf{k}_{q}\in\mathbb{R}^{r\times r}$ for $q=1,\dots,Q$. Therefore, for each new parameter $\bm{\theta}$ the construction of the reduced stiffness matrix $\mathbf{K}_{r}(\bm{\theta})$ requires the assembly of the full order stiffness matrix $\mathbf{K}(\bm{\theta})$ and its projection to the reduced space, so the reduced model remains coupled to the high dimension $N$. This bottleneck for multi-query scenarios is well-known such that hyper reduction techniques are developed like matrix gappy POD [matrixgappypod] or matrix DEIM [mdeim].
@@ 3.4 The operator inference approach for second order mechanical systems
In this subsection, we recall the main ideas of the operator inference approach applied to the linear second order mechanical system and closely follow [Peherstorfer2016Opinf, Filanova2023]. First, we write the reduced order system (23) (but for the case of input signal at the right hand side) for all time snapshots at once as follows:
$\mathbf{M}_{r}\ddot{\mathbf{U}}_{r}+\mathbf{K}_{r}(\bm{\theta})\mathbf{U_{r}}=\mathbf{B}_{r}\mathbf{Z},$ (25)
@@
$\mathbf{U}_{r}=\begin{bmatrix}|&\cdots&|\\ \mathbf{u}_{r}(\bm{\theta},t_{1})&\cdots&\mathbf{u}_{r}(\bm{\theta},t_{M})\\ |&\cdots&|\end{bmatrix}\in\mathbb{R}^{r\times M},\quad\ddot{\mathbf{U}}_{r}$ $=\begin{bmatrix}|&\cdots&|\\ \ddot{\mathbf{u}}_{r}(\bm{\theta},t_{1})&\cdots&\ddot{\mathbf{u}}_{r}(\bm{\theta},t_{M})\\ |&\cdots&|\end{bmatrix}\in\mathbb{R}^{r\times M}.$ (26)
@@
$\ddot{\mathbf{U}}_{r}+\mathbf{M}_{r}^{-1}\mathbf{K}_{r}(\bm{\theta})\mathbf{U_{r}}=\mathbf{M}_{r}^{-1}\mathbf{B}_{r}\mathbf{Z}.$ (27)
@@
where for a matrix $\mathbf{A}\in\mathbb{R}^{n\times m}$ the Frobenius norm is defined by $\left\lVert\mathbf{A}\right\rVert_{F}=\sqrt{\sum^{m}_{i=1}\sum^{n}_{j=1}{|a_{ij}|}^{2}}$ with $a_{ij}$ denoting the entries of $\mathbf{A}$.
The resulted operators $\widehat{\mathbf{K}}_{M}$ and $\widehat{\mathbf{B}}_{M}$ can then be used to integrate the second order mechanical system in (28) and obtain the operator inference reduced solution. Then by multiplying the reduced solution from the left by the matrix of POD modes $\mathbf{\Phi}$, one can approximate the FOM displacement. We note that this approach requires no access to the full order model code or algorithm. Therefore, it falls under the category of non-intrusive reduced order models or data-driven reduced order models. The approach so far assumes partial knowledge about the external forces represented in the input signals $\mathbf{z}$ and is termed as unconstrained operator inference [Filanova2023]. In the remaining part of this section we will recall another version of operator inference which is based on full knowledge of the forces. This version is termed forces-informed operator inference [Filanova2023].
@@
$\widehat{\mathbf{M}}=\mathbf{L}^{T}\mathbf{L},\quad\widehat{\mathbf{K}}=\mathbf{W}^{T}\mathbf{W},$ (35)
@@
$\min_{\mathbf{L},\mathbf{W}}\left\lVert\mathbf{L}^{T}\mathbf{L}\ddot{\mathbf{\widehat{U}}}+\mathbf{W}^{T}\mathbf{W}\mathbf{\widehat{U}}-\mathbf{\widehat{F}}\right\rVert_{F}^{2}.$ (36)
@@
We note that we consider operator inference here only for the linear mechanical model of Section 2.2 due to the complex structure of the nonlinear model. There exist approaches for operator inference for nonlinear models, see e.g. [PEHERSTORFER2016196, KPW24], where the nonlinearity requires to be of polynomial structure or can be lifted to polynomial structure using a variable transformation, see [Qian2020]. In [BENNER2020113433] an operator inference approach is introduced for systems with non-polynomial nonlinear terms which requires non-polynomial terms analytically.
@@ 3.5 Scaling the data for the operator inference approach
After solving (40), an approximation of the full displacement can be computed by $\mathbf{u}^{\text{OpInf}}(t)=\mathbf{\Phi}\mathbf{\widehat{u}}(t)\in\mathbb{R}^{N}$.

### v2

@@ 3.2 The intrusive POD-Galerkin ROM for second order mechanical systems
In this subsection, we describe the intrusive POD-Galerkin ROM for a mechanical system of second order without damping. For the derivation of a reduced order approximation of the full order model (FOM) (9) and (10), respectively, the FOM displacement field $\mathbf{u}(t)$ is approximated using a Galerkin ansatz as follows:
$\mathbf{u}(t)\approx\mathbf{\Phi}\mathbf{u}_{r}(t),$ (13)
with coefficient vector $\mathbf{u}_{r}(t)\in\mathbb{R}^{r}$ of dimension $r\ll N$ and reduced basis matrix $\mathbf{\Phi}=[\bm{\phi}_{1},\dots,\bm{\phi}_{r}]\in\mathbb{R}^{N\times r}$. Here, the reduced basis vectors $\{\bm{\phi}_{i}\}_{i=1}^{r}$ are generated using the POD method. For this a SVD of the snapshot matrix of the displacement can be used:
@@
By inserting the reduction assumption (13) in (9) and (10), respectively and performing a Galerkin projection onto the POD space, one may obtain the following linear reduced order dynamical system:
$\mathbf{\Phi}^{T}\mathbf{M}\mathbf{\Phi}\;\ddot{\mathbf{u}}_{r}(t)+\mathbf{\Phi}^{T}\mathbf{K}\mathbf{\Phi}\;\mathbf{u}_{r}(t)=\mathbf{\Phi}^{T}\mathbf{f}(t),\quad\textrm{for}\quad t\in[0,T],$ (15)
@@
$\mathbf{\Phi}^{T}\mathbf{M}\mathbf{\Phi}\;\ddot{\mathbf{u}}_{r}(t)+\mathbf{g}(\mathbf{\Phi}\mathbf{u}_{r}(t))=\mathbf{\Phi}^{T}\mathbf{f}(t),\quad\textrm{for}\quad t\in[0,T].$ (16)
@@
$\mathbf{M}_{r}=\mathbf{\Phi}^{T}\mathbf{M}\mathbf{\Phi},\quad\mathbf{K}_{r}=\mathbf{\Phi}^{T}\mathbf{K}\mathbf{\Phi},\quad\mathbf{f}_{r}(t)=\mathbf{\Phi}^{T}\mathbf{f}(t).$ (17)
@@
$\mathbf{M}_{r}\ddot{\mathbf{u}}_{r}(t)+\mathbf{K}_{r}\mathbf{u}_{r}(t)=\mathbf{f}_{r}(t),\quad\textrm{for}\quad t\in[0,T].$ (18)
@@ 3.3 The operator inference approach: motivation and data preparation
The basic idea of operator inference is to recover a numerical model directly from snapshot data by approximating the system operators. Usually, the structure of the system that resulted in the numerical data is known to a certain extent. This partial knowledge is used to formulate a mathematical problem for the approximation of the system operators. Since the estimation of the full order operators would be computationally immense, one directly estimates reduced system operators which are of much smaller dimension. In [63], this approach has been proposed and is termed as operator inference. The operator inference approach has also been extended to specific physical problems and to non-polynomial nonlinear cases specific physical problems, see, e.g. [9, 68, 84]. An operator inference method tailored for second order mechanical systems has been proposed in [8, 26], which we closely follow here to apply it to our mechanical systems of interest.
In the operator inference approach, derivative information of the solution data is required. In many occasions, the numerical solver may provide this data. If that is not the case, then one may approximate the derivatives using finite difference schemes, e.g. using an eighth order finite difference scheme, see [77]. The second derivatives of the displacement are assembled in the following snapshot matrix:
$\ddot{\mathbf{U}}$ $=\begin{bmatrix}|&\cdots&|\\ \ddot{\mathbf{u}}(t_{1})&\cdots&\ddot{\mathbf{u}}(t_{M})\\ |&\cdots&|\end{bmatrix}\in\mathbb{R}^{N\times M}.$ (20)
$\mathbf{\widehat{U}}=\mathbf{\Phi}^{T}\mathbf{U},\quad{\mathbf{\widehat{\ddot{U}}}}=\mathbf{\Phi}^{T}\ddot{\mathbf{U}},$ (21)
where $\mathbf{\Phi}$ is the matrix of POD modes computed using e.g. a singular value decomposition (SVD) of $\mathbf{U}$, see Section 3.2.
@@ 3.4 The operator inference approach for second order mechanical systems
In this subsection, we recall the main ideas of the operator inference approach applied to the linear second order mechanical system and closely follow [63, 26]. First, we write the reduced order system (18) (but for the case of input signal (19) at the right hand side) for all time snapshots at once as follows:
$\mathbf{M}_{r}\ddot{\mathbf{U}}_{r}+\mathbf{K}_{r}\mathbf{U_{r}}=\mathbf{B}_{r}\mathbf{Z},$ (22)
@@
$\mathbf{U}_{r}=\begin{bmatrix}|&\cdots&|\\ \mathbf{u}_{r}(t_{1})&\cdots&\mathbf{u}_{r}(t_{M})\\ |&\cdots&|\end{bmatrix}\in\mathbb{R}^{r\times M},\quad\ddot{\mathbf{U}}_{r}$ $=\begin{bmatrix}|&\cdots&|\\ \ddot{\mathbf{u}}_{r}(t_{1})&\cdots&\ddot{\mathbf{u}}_{r}(t_{M})\\ |&\cdots&|\end{bmatrix}\in\mathbb{R}^{r\times M}.$ (23)
@@
$\ddot{\mathbf{U}}_{r}+\mathbf{M}_{r}^{-1}\mathbf{K}_{r}\mathbf{U_{r}}=\mathbf{M}_{r}^{-1}\mathbf{B}_{r}\mathbf{Z}.$ (24)
@@
where for a matrix $\mathbf{A}\in\mathbb{R}^{n\times m}$ the Frobenius norm is defined by $\left\lVert\mathbf{A}\right\rVert_{F}=\sqrt{\sum^{n}_{i=1}\sum^{m}_{j=1}{|a_{ij}|}^{2}}$ with $a_{ij}$ denoting the entries of $\mathbf{A}$.
The optimization problem (26) can be written in a more compact way by combining the data matrices as follows
$\widehat{\mathbf{D}}=\begin{bmatrix}\mathbf{\widehat{U}}\\ \mathbf{Z}\end{bmatrix}$, (27)
and by defining the following combined unknown operators:
$\widehat{\mathbf{P}}=\left[-\widehat{\mathbf{K}}_{M},\,\widehat{\mathbf{B}}_{M}\right],$ (28)
which allows to write the optimization problem as follows:
$\min_{\widehat{\mathbf{P}}\in\mathbb{R}^{r\times(r+N_{I})}}\lVert\widehat{\mathbf{P}}\widehat{\mathbf{D}}-\ddot{\widehat{\mathbf{U}}}\rVert_{F}^{2}.$ (29)
The solution to the minimization problem above satisfies the normal equations:
$\widehat{\mathbf{P}}\widehat{\mathbf{D}}\widehat{\mathbf{D}}^{T}=\ddot{\widehat{\mathbf{U}}}\widehat{\mathbf{D}}^{T}.$ (30)
The equation above has a unique solution if and only if $\widehat{\mathbf{D}}^{T}\widehat{\mathbf{D}}$ is invertible which is fulfilled when $\widehat{\mathbf{D}}$ has full rank. The resulted operator $\widehat{\mathbf{P}}$ can then be used to integrate the second order mechanical system in (25) and obtain the operator inference reduced solution. Then by multiplying the reduced solution from the left by the matrix of POD modes $\mathbf{\Phi}$, one can approximate the FOM displacement. We note that this approach requires no access to the full order model code or algorithm. Therefore, it falls under the category of non-intrusive reduced order models or data-driven reduced order models. The approach so far assumes partial knowledge about the external forces represented in the input signals $\mathbf{z}$ and is termed as unconstrained operator inference [26]. In the remaining part of this section we will recall another version of operator inference which is based on full knowledge of the forces. This version is termed forces-informed operator inference [26].
@@
In the current case, we define the combined data matrix:
$\widehat{\mathbf{D}}=\begin{bmatrix}\ddot{\mathbf{\widehat{U}}}\\ \mathbf{\widehat{U}}\end{bmatrix}$, (36)
Then, the operator inference forces-informed problem is given by:
$\min_{\widehat{\mathbf{M}}\succ 0,\widehat{\mathbf{K}}\succ 0}\left\lVert\left[\widehat{\mathbf{M}}\quad\widehat{\mathbf{K}}\right]\widehat{\mathbf{D}}-\mathbf{\widehat{F}}\right\rVert_{F}^{2}.$ (37)
@@
$\widehat{\mathbf{M}}=\mathbf{L}^{T}\mathbf{L},\quad\widehat{\mathbf{K}}=\mathbf{E}^{T}\mathbf{E},$ (38)
@@
$\min_{\mathbf{L},\mathbf{E}}\left\lVert\mathbf{L}^{T}\mathbf{L}\ddot{\mathbf{\widehat{U}}}+\mathbf{E}^{T}\mathbf{E}\mathbf{\widehat{U}}-\mathbf{\widehat{F}}\right\rVert_{F}^{2}.$ (39)
@@
We note that we consider operator inference here only for the linear mechanical model of Section 2.2 due to the complex structure of the nonlinear model. In previous work [68], the nonlinearity requires to be of polynomial structure or can be lifted to polynomial structure using a variable transformation. In [9] an operator inference approach is introduced for systems with non-polynomial nonlinear terms which requires non-polynomial terms analytically.
@@ 3.5 Scaling the data for the operator inference approach
After solving (43), an approximation of the full displacement can be computed by multiplying the reduced solution from the left by the matrix of POD modes $\mathbf{\Phi}$.
3.6 The discrete operator inference approach for second order mechanical systems
In Section 3.4, we have presented the operator inference approach for the second order mechanical system of interest written in a semi-discretized setting. However, the operator inference method could be also applied in a fully discretized setup. We refer the reader to the works [24, 64] for an application of discrete OpInf to first order systems.
The continuous time formulation of OpInf requires the evaluation of time derivatives of the displacement field. These derivative data are typically obtained through finite difference approximations when they are not provided directly by the numerical solver. In practice, however, estimating derivatives from data may introduce significant numerical errors, particularly when the data are noisy or when the temporal resolution is limited. Such errors propagate directly into the regression problem used to identify the reduced operators and may deteriorate the stability and predictive accuracy of the inferred model.
An alternative strategy consists in formulating the operator inference problem directly at the level of the time discretized dynamics. Depending on the time integration scheme used for the full order model, the discrete evolution of the displacement field can be expressed as a recurrence relation linking several consecutive time levels. For instance, explicit or semi-implicit schemes such as central-difference or theta-type methods lead to relations of the following form:
$\mathbf{G}\mathbf{u}_{k+2}=\mathbf{H}\mathbf{u}_{k+1}+\mathbf{R}\mathbf{u}_{k}+\mathbf{C}\mathbf{z}_{k+2},$ (44)
where $\mathbf{G}\in\mathbb{R}^{N\times N}$, $\mathbf{H}\in\mathbb{R}^{N\times N}$, $\mathbf{R}\in\mathbb{R}^{N\times N}$ and $\mathbf{C}\in\mathbb{R}^{N\times N_{I}}$. Furthermore, $\mathbf{u}_{k}=\mathbf{u}(t_{k})$ denotes the displacement vector at time $t_{k}$, while $\mathbf{z}_{k}=\mathbf{z}(t_{k})$ represents the input signal. Equation (44) thus defines an update relation for computing the displacement vector $\mathbf{u}_{k+2}$ based on the two previous displacement states $\mathbf{u}_{k}$, $\mathbf{u}_{k+1}$ and the input vector $\mathbf{z}_{k}$.
It should be noted that not all full order time integration schemes can be written in the form of (44). For instance, certain methods (such as the Newmark scheme) evolve the displacement and velocity variables simultaneously and therefore do not admit a recurrence relation involving displacement variables alone. Consequently, the fully discrete formulation of OpInf should be viewed as an additional modeling option that may be advantageous when the underlying time integration scheme naturally leads to a displacement-based recurrence relation.
Similar to (13), a Galerkin ansatz could be introduced for the displacement vector $\mathbf{u}_{k}$ as follows:
$\mathbf{u}_{k}\approx\mathbf{\Phi}\mathbf{u}_{k,r},$ (45)
where the reduced displacement at time step $t_{k}$ is $\mathbf{u}_{k,r}\in\mathbb{R}^{r}$. The substitution of (45) into (44) and the projection onto the reduced space spanned by the columns of $\mathbf{\Phi}$ yields:
$\mathbf{G}_{r}\mathbf{u}_{r,k+2}=\mathbf{H}_{r}\mathbf{u}_{r,k+1}+\mathbf{R}_{r}\mathbf{u}_{r,k}+\mathbf{C}_{r}\mathbf{z}_{k+2},$ (46)
where
$\mathbf{G}_{r}=\mathbf{\Phi}^{T}\mathbf{G}\mathbf{\Phi},\quad\mathbf{H}_{r}=\mathbf{\Phi}^{T}\mathbf{H}\mathbf{\Phi},\quad\mathbf{R}_{r}=\mathbf{\Phi}^{T}\mathbf{R}\mathbf{\Phi},\quad\mathbf{C}_{r}=\mathbf{\Phi}^{T}\mathbf{C}.$ (47)
The solution of the FOM and the ROM in (44) and (46), respectively, is carried out by inverting the matrix $\mathbf{G}$ or $\mathbf{G}_{r}$ (assuming that they are invertible). In this section, we assume $\mathbf{G}_{r}=\mathbf{I}$. From the perspective of operator inference, this assumption does not restrict generality since the discrete system can be left-multiplied by $\mathbf{G}_{r}^{-1}$. Consequently, the operators identified by the inference procedure correspond to the effective matrices $\mathbf{G}_{r}^{-1}\mathbf{H}_{r}$, $\mathbf{G}_{r}^{-1}\mathbf{R}_{r}$, and $\mathbf{G}_{r}^{-1}\mathbf{C}_{r}$.
Unlike the time continuous formulation, the operator inference approach developed for the discrete time setting does not require time derivative data. The first step of the discrete OpInf approach consists of decomposing the displacement data into three matrices: the snapshot data, the one-step–shifted snapshots, and the two-step–shifted snapshots, defined as follows:
$\mathbf{U}_{0}$ $=\begin{bmatrix}|&\cdots&|\\ \mathbf{u}_{1}&\cdots&\mathbf{u}_{M-2}\\ |&\cdots&|\end{bmatrix}\in\mathbb{R}^{N\times M-2},\quad\mathbf{U}_{1}=\begin{bmatrix}|&\cdots&|\\ \mathbf{u}_{2}&\cdots&\mathbf{u}_{M-1}\\ |&\cdots&|\end{bmatrix}\in\mathbb{R}^{N\times M-2},\quad$ (48)
$\mathbf{U}_{2}$ $=\begin{bmatrix}|&\cdots&|\\ \mathbf{u}_{3}&\cdots&\mathbf{u}_{M}\\ |&\cdots&|\end{bmatrix}\in\mathbb{R}^{N\times M-2},\quad\mathbf{Z}_{2}=\begin{bmatrix}|&\cdots&|\\ \mathbf{z}_{3}&\cdots&\mathbf{z}_{M}\\ |&\cdots&|\end{bmatrix}\in\mathbb{R}^{N_{I}\times M-2}.$ (49)
The corresponding reduced data is computed as follows
$\mathbf{\widehat{U}}_{0}=\mathbf{\Phi}^{T}\mathbf{U}_{0},\quad\mathbf{\widehat{U}}_{1}=\mathbf{\Phi}^{T}\mathbf{U}_{1},\quad\mathbf{\widehat{U}}_{2}=\mathbf{\Phi}^{T}\mathbf{U}_{2}.$ (50)
Then, the least squares regression problem for operator inference reads as follows:
$\min_{\widehat{\mathbf{P}}\in\mathbb{R}^{r\times(2r+N_{I})}}\lVert\widehat{\mathbf{P}}\widehat{\mathbf{D}}-\mathbf{\widehat{U}}_{2}\rVert_{F}^{2},$ (51)
where
$\widehat{\mathbf{D}}=\begin{bmatrix}\mathbf{\widehat{U}}_{1}\\ \mathbf{\widehat{U}}_{0}\\ \mathbf{Z}_{2}\end{bmatrix},\quad\widehat{\mathbf{P}}=\left[\widehat{\mathbf{H}},\,\widehat{\mathbf{R}},\,\widehat{\mathbf{C}}\right$]. (52)
3.7 The re-projection method for second order mechanical systems
In this section we propose an extension of the re-projection method presented in [64] to the case of second order mechanical systems. The re-projection method aims at generating reduced trajectories (which we denote with a bar $\mathbf{\bar{U}}$) which are identical to the ones produced by the intrusive reduced operators. This results in a zero closure error $\lVert\mathbf{U}_{r}-\mathbf{\bar{U}}\rVert_{F}=0$. It has been shown in [64] that OpInf applied on the re-projected trajectories recovers the intrusive reduced operators under certain conditions.
The re-projection method aims at generating snapshots trajectories which enforce a zero closure error. The closure error is caused by the non-Markovian dynamics of the projected trajectories. The non-Markovian dynamics are not produced by the reduced order model and by eliminating them one could recover the trajectories generated by the reduced order model. This is done by a data sampling procedure which cancels out the non-Markovian dynamics in the reduced space. We introduce the re-projection method in Algorithm 1 and Algorithm 2 for the time continuous case and the time discrete case, respectively. It can be seen from Algorithm 1 that we need to have both the first and second time derivative of the full order solve at each iteration in order to perform the method. The OpInf problem in the continuous time case will read as follows:
$\min_{\bar{\mathbf{P}}\in\mathbb{R}^{r\times(r+N_{I})}}\lVert\bar{\mathbf{P}}\bar{\mathbf{D}}-\mathbf{\ddot{\bar{U}}}\rVert_{F}^{2},\quad\text{where}\quad\bar{\mathbf{D}}=\begin{bmatrix}\mathbf{\bar{U}}\\ \mathbf{Z}\end{bmatrix},\quad\text{and}\quad\bar{\mathbf{P}}=\left[-\bar{\mathbf{K}}_{M},\,\bar{\mathbf{B}}_{M}\right].$ (53)
On the other hand, the discrete OpInf problem is:
$\min_{\bar{\mathbf{P}}\in\mathbb{R}^{r\times(2r+N_{I})}}\lVert\bar{\mathbf{P}}\bar{\mathbf{D}}-\mathbf{\bar{U}}_{2}\rVert_{F}^{2},\quad\text{where}\quad\bar{\mathbf{D}}=\begin{bmatrix}\mathbf{\bar{U}}_{1}\\ \mathbf{\bar{U}}_{0}\\ \mathbf{Z}_{2}\end{bmatrix},\quad\text{and}\quad\bar{\mathbf{P}}=\left[\bar{\mathbf{H}},\,\bar{\mathbf{R}},\,\bar{\mathbf{C}}\right].$ (54)
Algorithm 1 The re-projection method for second order mechanical system (time continuous case)
1: Start with the initial displacement $\mathbf{u}_{1}$ and the initial time derivative of the displacement $\mathbf{\dot{u}}_{1}$ with the input signal $\mathbf{z}_{1}$.
2: Solve the FOM to obtain $\mathbf{\ddot{u}}_{1}$.
3: Compute $\mathbf{\bar{u}}_{1}=\mathbf{\Phi}^{T}\mathbf{u}_{1}\,\mathbf{\dot{\bar{u}}}_{1}=\mathbf{\Phi}^{T}\mathbf{\dot{u}}_{1}\,\mathbf{\ddot{\bar{u}}}_{1}=\mathbf{\Phi}^{T}\mathbf{\ddot{u}}_{1}$
4:for $i=2,\ldots,M$ do
5:  Solve the FOM for one time step using $\mathbf{\Phi}\mathbf{\bar{u}}_{i-1},\,\mathbf{\Phi}\mathbf{\dot{\bar{u}}}_{i-1},\,\mathbf{\Phi}\mathbf{\ddot{\bar{u}}}_{i-1},\,\mathbf{z}_{i-1}\rightarrow\mathbf{u}_{i}^{\star},\,\mathbf{\dot{u}}_{i}^{\star},\,\mathbf{\ddot{u}}_{i}^{\star}$.
6:  Set $\mathbf{\bar{u}}_{i}=\mathbf{\Phi}^{T}\mathbf{u}_{i}^{\star},\,\mathbf{\dot{\bar{u}}}_{i}=\mathbf{\Phi}^{T}\mathbf{\dot{u}}_{1}^{\star},\,\mathbf{\ddot{\bar{u}}}_{i}=\mathbf{\Phi}^{T}\mathbf{\ddot{u}}_{i}^{\star}$.
7:end for
8: Return the re-projected reduced trajectories $\mathbf{\bar{U}}=[\mathbf{\bar{u}}_{1},\mathbf{\bar{u}}_{2},\ldots,\mathbf{\bar{u}}_{M}]$ and $\mathbf{\ddot{\bar{U}}}=[\mathbf{\ddot{\bar{u}}}_{1},\mathbf{\ddot{\bar{u}}}_{2},\ldots,\mathbf{\ddot{\bar{u}}}_{M}]$
Algorithm 2 The re-projection method for second order mechanical system (time discrete case)
1: Start with the displacement at the first two steps $\mathbf{u}_{1}$ and $\mathbf{u}_{2}$ with the input signal discretized vector $\mathbf{z}_{1}$.
2: Compute $\mathbf{\bar{u}}_{1}=\mathbf{\Phi}^{T}\mathbf{u}_{1}\,\mathbf{\bar{u}}_{2}=\mathbf{\Phi}^{T}\mathbf{u}_{2}$.
3:for $i=2,\ldots,M-1$ do
4:  Solve the FOM for one time step using $\mathbf{\Phi}\mathbf{\bar{u}}_{i-1},\,\mathbf{\Phi}\mathbf{\bar{u}}_{i},\,\mathbf{z}_{i+1}\rightarrow\mathbf{u}_{i+1}^{\star}.$
5:  Set $\mathbf{\bar{u}}_{i+1}=\mathbf{\Phi}^{T}\mathbf{u}_{i+1}^{\star}$.
6:end for
7: Return the re-projected reduced trajectories $\mathbf{\bar{U}}_{0}=[\mathbf{\bar{u}}_{1},\mathbf{\bar{u}}_{2},\ldots,\mathbf{\bar{u}}_{M-2}]$, $\mathbf{\bar{U}}_{1}=[\mathbf{\bar{u}}_{2},\mathbf{\bar{u}}_{3},\ldots,\mathbf{\bar{u}}_{M-1}]$ and $\mathbf{\bar{U}}_{2}=[\mathbf{\bar{u}}_{3},\mathbf{\bar{u}}_{4},\ldots,\mathbf{\bar{u}}_{M}]$.
As already mentioned in [62, 64], we have observed that in several cases one may need to augment the data matrix $\bar{\mathbf{D}}$ or $\widehat{\mathbf{D}}$ with trajectories coming from different input signals in order at the end to reduce the condition number of the matrix $\bar{\mathbf{D}}\bar{\mathbf{D}}^{T}$ or $\widehat{\mathbf{D}}\widehat{\mathbf{D}}^{T}$ which could pollute the learned operators and introduce numerical errors. For instance in the time continuous case, we assume that we have $q$ different input signals ${[\mathbf{z}^{j}(t)]}_{j=1}^{q}$ which give the matrices ${[\mathbf{Z}^{j}]}_{j=1}^{q}$, then we apply Algorithm 1 $q$ times to obtain $\mathbf{\bar{U}^{1}},\dots,\mathbf{\bar{U}^{q}}$ and $\mathbf{\ddot{\bar{U}}^{1}},\dots,\mathbf{\ddot{\bar{U}}^{q}}$. We concatenate the matrices to obtain:
$\mathbf{\bar{U}}=[\mathbf{\bar{U}^{1}},\mathbf{\bar{U}^{2}},\ldots,\mathbf{\bar{U}^{q}}],\quad\mathbf{\ddot{\bar{U}}}=[\mathbf{\ddot{\bar{U}}^{1}},\mathbf{\ddot{\bar{U}}^{2}},\ldots,\mathbf{\ddot{\bar{U}}^{q}}],\quad\mathbf{\bar{Z}}=[\mathbf{Z^{1}},\mathbf{Z^{2}},\ldots,\mathbf{Z^{q}}],$ (55)
then we solve the OpInf (53) with the matrices above.